OpenStudy (anonymous):

use implicit differentiation to find the slope of the tangent line at (-2,3) and use it to find the equation of the tangent line in y=mx + b form x^2+y^2=13

OpenStudy (experimentx):

differentiate it 2x+2y*dy/dx = 0; or dy/dx = -x/y put the value m = 2/3 since you know point, find the equation

OpenStudy (anonymous):

ok ya thats what i got for the slope but i don't know how to find b

OpenStudy (experimentx):

put the values -2, 3 on that equation and value of slope, you will find b

OpenStudy (anonymous):

ok well i got y=2/3x+5/3 but thats not right either..

OpenStudy (experimentx):

b = 3 + 2*4/3

OpenStudy (experimentx):

sorry, 3+2*2/3